OpenStudy (anonymous):

find the maximum y value on the graph of y=f(x)...f(x)=x^2+4x+10

OpenStudy (campbell_st):

given the curve is concave down the maximum y values occur at (-infinity, infinity) perhaps you have mistyped...

OpenStudy (anonymous):

sorry it's -x^2+4x+10

OpenStudy (campbell_st):

ok.... that makes nore sense find the line of symmetry using \[x = \frac{-b}{2a}\] then substitute it into the equation and evaluate. Another method is to complete the square and write your equation in vertex form \[y = (x -h)^2 + k \] for me the line of symmetry and substitution is easier.

OpenStudy (campbell_st):

the general for is y = ax^2 + bx + c in your question a = -1 and b = 4 the formula is the 1st part of the general quadratic formula. hope it helps

OpenStudy (anonymous):

@campbell is the answer -14

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

isn't it -(2)^2 + 4(2) + 10 = 14